• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name efuchey

    Log entry time 18:41:31 on September 27, 2010

    Entry number 332971


    Summary of compton status: - I did manage to lock the cavity this morning, did not manage this afternoon though; - I was able to setup the DAQ; - finally, I managed to take LED data (attached figure).

    I shall find out what worked this morning which didnt work this afternoon.

    A copy of this log entry has been emailed to: brads@jlab.org, mfriend@jlab.org, dvcs@jlab.org

    Figure 1